# 2cm2cm1.5cm1.5cm

### v1

@@ 3.2. Rotations
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ such that the origin is in the interior of $L\cap K,$ and there are no
  1. (1)
    contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L$ and $K;$
  2. (2)
    positive weights $\upalpha_{1},\dots,\upalpha_{m}$
satisfying identity (1.5). Then for some $\varepsilon>0$, there is a differentiable path $R_{t},$ $t\in[0,\varepsilon]$ of orthogonal transofrmations such that $R_{t}K\subset\mathrm{int}L$ for all $t\in(0,\varepsilon].$
@@
There is nothing to prove if the set of contact pairs is empty. We assume that the cone $\operatorname{Cone}_{\mathrm{op}}=\mathrm{Pos}\ \!{\left\{v\otimes p:\;(v,p)\text{ is a contact pair of }K\text{ and }L\right\}}$ is non-empty.
@@
$\mathcal{M}_{\mathrm{Sym}}^{d}\cap\operatorname{Cone}_{\mathrm{op}}=0.$
Thus, there is non-zero $W$ separating the two convex cones. However, $\mathcal{M}_{\mathrm{Sym}}^{d}$ is a subspace, hence $W\in{\left(\mathcal{M}_{\mathrm{Sym}}^{d}\right)}^{\bot},$ that is, $W$ is skew symmetric.
Set $R_{t}=e^{tW},$ $H=0,$ and ${U}^{\top}=\mathrm{Id}_{d}$ in (2.1). Clearly, $R_{t},t\in[0,1]$ is a differentiable path of orthogonal matrices in $\mathcal{M}^{d}$ satisfying (2.2). The result follows from Corollary 2.7. ∎
@@ 3.3. Search for a good center
We show how $z$ can be found to obtain (1.3) in Theorems 1.6 and 1.4.
@@
satisfying identities (1.1) – (1.4). Then there exists a point $z$ in the interior of $L\cap K$, contact pairs $\{(v_{1}^{\prime},p_{1}^{\prime}),\dots,(v_{m}^{\prime},p_{m}^{\prime})\}$ of $K-z$ and $L-z$, and weights $c_{1},\dots,c_{m}>0$ such that
${\sum_{i\in[m]}c_{i}v_{i}^{\prime}\otimes p_{i}^{\prime}}=\sum_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i},\quad\text{and}\quad\sum_{i\in[m]}c_{i}P_{F}v_{i}^{\prime}=\sum_{i\in[m]}c_{i}p_{i}^{\prime}=\mathbf{0}.$
@@
This is a purely linear algebraic observation.
We aim at finding a point $z$ in the interior of $K$ such that when we shift the origin to $z$, the shifted contact pairs $(v_{i}-z,p_{i}^{\prime})$ satisfy the required identities for $K-z$ and $L-z$.
First, observe how a contact pair transforms under translation. A pair $(v,p)$ is a contact pair of $K$ and $L$ if and only if $v$ belongs to the intersection of the boundaries of $K$ and $L$, and the half-space
@@
Set $M=\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}.$ Next, we verify the identity $\sum\limits_{i\in[m]}c_{i}(z)v_{i}(z)\otimes p_{i}(z)=M$. Indeed,
$\sum_{i\in[m]}c_{i}(z)v_{i}(z)\otimes p_{i}(z)=\sum_{i\in[m]}\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)(v_{i}-z)\otimes\frac{p_{i}}{1-\left\langle z,p_{i}\right\rangle}=\sum_{i\in[m]}\upalpha_{i}(v_{i}-z)\otimes p_{i}.$
Simplifying, we get
@@
$\sum_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)$ $=\sum_{i\in[m]}\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)P_{F}(v_{i}-z)$
@@
Using the identities $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and $P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)P_{F}=P_{F}$, we can simplify:
$\sum\limits_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)=\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}-P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)P_{F}z-P_{F}z\sum\limits_{i\in[m]}\upalpha_{i}+P_{F}z\left\langle z,\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\right\rangle$
$=\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}-P_{F}z-P_{F}z\left(\sum\limits_{i\in[m]}\upalpha_{i}\right)+P_{F}z\cdot 0=\left(\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}\right)-(d+1)P_{F}z.$
Thus, setting
$z=\frac{\sum\limits_{i\in[m]}\upalpha_{i}v_{i}}{d+1}$
ensures that $\sum\limits_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)=\mathbf{0}$, as desired. Since $\sum\limits_{i\in[m]}\upalpha_{i}=d$, and $\mathbf{0}\in\mathrm{int}K$, we have $z\in\mathrm{int}K$. ∎
Proof of Theorem 1.6.
Combining Theorem 3.2 with Lemma 3.5 yields our theorem. ∎
Proof of Theorem 1.4.
Combining Theorem 3.3 with Lemma 3.5 yields our theorem. ∎
3.4. The condition is sufficient for local extremality – in some cases
the positive $F$-positions of $S$.
Theorem 3.6.
4. Containment
Let $K_{\text{in}}$ and $K_{\text{out}}$ be convex bodies in $\mathbb{R}^{d}$ such that $K_{\text{in}}\subset K_{\text{out}}$ and $K_{\text{in}}$ contains the origin in its interior.
Assume there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K_{\text{in}}$ and $K_{\text{out}},$ and weights $\upalpha_{1},\dots,\upalpha_{m}$ such that identities (1.1) – (1.4) hold.
Denote $L_{\text{in}}=\operatorname{conv}\left\{v_{1},\dots,v_{m}\right\}$ and $L_{\text{out}}=\bigcap\limits_{i\in[m]}\left\{x\in\mathbb{R}^{d}:\;\left\langle p_{i},x\right\rangle\leq 1\right\}.$
We call $L_{\text{in}}$ and $L_{\text{out}}$ the inner and outer contact polytope corresponding to the contact pairs $(v_{i},p_{i})$. Clearly,
$L_{\text{in}}\subset K_{\text{in}}\subset K_{\text{out}}\subset L_{\text{out}}.$
Lemma 4.1.
Under the above notation, we have
$L_{\mathrm{out}}\cap F\subset-d\cdot P_{F}L_{\mathrm{in}}.$
Proof.
@@ 4. Containment
$\sum\limits_{i\in[m]}\upalpha_{i}(1-\left\langle p_{i},x\right\rangle)P_{F}v_{i}\in\sum\limits_{i\in[m]}\upalpha_{i}\cdot P_{F}L_{\text{in}}\cdot(1-\left\langle p_{i},x\right\rangle)\stackrel{{\scriptstyle\left\langle p_{i},x\right\rangle\leq 1}}{{=}}$
@@
Clearly, this lemma yields Corollary 1.5 as well as (3) in Theorem 1.7.
@@ 4.1. No containment in the whole space
Contrasting Corollary 1.5, we show that there is no meaningful inclusion in ${F}^{\bot}$. Surprisingly, the example is provided by ellipsoids.
Lemma 4.2.
Fix $20\leq s\leq d-$ 2 and a positive number $\lambda$. Let $E$ be an origin-centered ellipsoid in $\mathbb{R}^{d}$ whose principal semi-axes have lengths $\lambda,\ldots,\lambda^{d}$. Let $E^{\prime}\in\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$ be the solution to Problem 2 for some fixed $s$-dimensional subspace $F$. Then
@@
Example 4.4.
Let $K$ be a convex body in $\mathbb{R}^{d}$, and let $A:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the diagonal linear operator with diagonal entries, $\lambda,\ldots,\lambda^{d}$. Let $K^{\prime}=A^{\prime}(K)$ be a solution to Problem 4 with some linear operator $A^{\prime}$, for $K^{\prime\prime}=A(K)$, with any $d-s\geq 2$. Then, for any $\mu>0$ there is some value of $\lambda>0$ such that
@@
    for any $\mu^{\prime}<\mu$, no homothetic copy $x+\mu^{\prime}K^{\prime}$ contains $K^{\prime\prime}$,
@@
    $\frac{\operatorname{vol}\nolimits_{d}K^{\prime}}{\operatorname{vol}\nolimits_{d}K^{\prime\prime}}\leq\frac{1}{\mu}$.
@@
Let $B_{\mathrm{is}}$ be a largest ball contained in $K$, and let $B_{\mathrm{cs}}$ be a smallest ball containing $K$. Assume that $B_{\mathrm{cs}}=y+\nu B_{\mathrm{is}}$ for $\nu\geq 1$, $y\in\mathbb{R}^{d}$.
Let $\bar{\mu}>0$ be the smallest number such that a homothetic copy of $K^{\prime}$ of ratio $\bar{\mu}$ contains $K^{\prime\prime}$. Then there is a homothetic copy of $A^{\prime}(B_{\mathrm{cs}})$ of ratio $\bar{\mu}$ that contains $A(B_{\mathrm{is}})$. Consequently, there is a homothetic copy of $A^{\prime}(B_{\mathrm{is}})$ of ratio $\bar{\mu}\nu^{2}$ that contains $A(B_{\mathrm{cs}})$. Clearly, $A^{\prime}(B_{\mathrm{is}})\subseteq A(B_{\mathrm{cs}})$. Thus, by the previous example, $\bar{\mu}\nu^{2}\geq\lambda$, implying that $\bar{\mu}\geq\frac{\lambda}{\bar{\mu}^{2}}$, proving (a).
@@ 5.1. Contact points on the sphere
Let the origin-centered ellipsoid $E=A\mathbf{B}^{d}$ be a solution to the $F$-revolution John problem for a convex body $K$ in $\mathbb{R}^{d}$. Then there are unit vectors $u_{1},\dots,u_{m}$, and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that $(Au_{1},{A}^{-1}u_{1}),\dots,(Au_{m},{A}^{-1}u_{m})$ are contact points of $E$ and $K$, and
@@
Note that a standard argument using Carathéodory’s theorem shows that the number $m$ of points needed in Corollary 5.1 is at most $12d^{2}+d+$ 1.
@@
Assume that $E=A\mathbf{B}^{d}$ is a solution to the $F$-revolution John problem with any axis for a convex body $K$ in $\mathbb{R}^{d}$. Then there exist $m\leq 2d^{2}+d+1$, unit vectors $u_{1},\dots,u_{m}$, and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that the conclusions of Corollary 5.1 hold as well as
@@
We will call a position $K^{\prime}\in\mathcal{A}_{F}\!\left(K\right)$ with fixed axis $F$ of a convex body $K$ the John position with fixed axis $F$, if the standard unit ball is the solution to John’s general problem with fixed axis $F$ for $K^{\prime}$ and $\mathbf{B}^{d}.$

### v2

@@ 3.2. Rotations
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ such that the origin is in the interior of $L\cap K,$ and there are no contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $L$ and $K$ and corresponding positive weights $\upalpha_{1},\dots,\upalpha_{m}$ satisfying identity (1.4). Then for some $\varepsilon>0$, there is a differentiable path $R_{t},$ $t\in(0,\varepsilon]$ of orthogonal transformations such that $R_{t}K\subset\mathrm{int}L$ for all $t\in(0,\varepsilon].$
@@
There is nothing to prove if the set of contact pairs is empty. We assume that the cone $\operatorname{Cone}_{\mathrm{op}}=\mathrm{Pos}\ \!{\left\{p\otimes v:\;(v,p)\text{ is a contact pair of }K\text{ and }L\right\}}$ is non-empty. Clearly, it is a closed set.
@@
$\mathcal{M}_{\mathrm{Sym}}^{d}\cap\operatorname{Cone}_{\mathrm{op}}=\left\{\mathbf{0}_{\mathcal{M}^{d}}\right\}.$
Thus, there is a non-zero matrix $W$ separating the closed convex cone $\operatorname{Cone}_{\mathrm{op}}$ from $\mathcal{M}_{\mathrm{Sym}}^{d}$ in the following sense: for every nonzero $M_{1}\in\mathcal{M}_{\mathrm{Sym}}^{d}$ and $M_{2}\in\operatorname{Cone}_{\mathrm{op}}$, we have $\left\langle M_{1},W\right\rangle_{\mathcal{M}^{d}}\leq 0<\left\langle M_{2},W\right\rangle_{\mathcal{M}^{d}}$. Since $\mathcal{M}_{\mathrm{Sym}}^{d}$ is a subspace, $W\in{\left(\mathcal{M}_{\mathrm{Sym}}^{d}\right)}^{\bot},$ that is, $W$ is skew-symmetric.
Set $R_{t}=e^{tW}$ and $H=0$ in (2.1). Clearly, $R_{t},t\in[0,1]$ is a differentiable path of orthogonal matrices in $\mathcal{M}^{d}$ satisfying (2.2). By direct computation,
$\left.\frac{\,\mathrm{d}}{\,\mathrm{d}t}\right|_{t=0^{+}}\left(R_{t}v\right)=Wv.$
The result follows from the descriptions of the tangent cone in Lemma 2.2 and Lemma 2.3, and standard compactness arguments.
3.3. The condition is sufficient for local extremality – in some cases
the positive $F$-positions of $S$.
Theorem 3.3.
We leave the proof to the reader.
4. Containment
Let us recall how the authors of [GLMP04] derived Theorem 1.5 from Theorem 1.4 in the case $F=\mathbb{R}^{d}$. Note that the identities (1.1), (1.2), and (1.3) in Theorem 1.4 can be rewritten with respect to an arbitrary point in the interior of $K$; in other words, we may choose the origin arbitrarily within the interior of $K$.
In [GLMP04] it is shown that in the case $F=\mathbb{R}^{d}$, one can choose the origin so that, in addition to the identities (1.1), (1.2), and (1.3), the following identity also holds
$\sum\limits_{i\in[m]}\upalpha_{i}v_{i}=0.$
As we will show, in our setting, the following relation suffices to prove the containment in Theorem 1.5:
(4.1) $\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}=\mathbf{0}.$
4.1. Inclusion for contact polytopes
Let $K$ and $L$ be convex bodies in $\mathbb{R}^{d}$ such that $K\subset L$ and $K$ contains the origin in its interior. Let $F$ be a non-trivial subspace of $\mathbb{R}^{d}.$
Assume that there are contact pairs $(v_{1},p_{1}),\dots,(v_{m},p_{m})$ of $K$ and $L,$ and weights $\upalpha_{1},\dots,\upalpha_{m}$ such that identities (1.1) – (1.3) hold, and moreover, identity (4.1) holds.
Denote $K_{\text{in}}=\operatorname{conv}\left\{v_{1},\dots,v_{m}\right\}$ and $L_{\text{out}}=\bigcap\limits_{i\in[m]}\left\{x\in\mathbb{R}^{d}:\;\left\langle p_{i},x\right\rangle\leq 1\right\}.$
We call $K_{\text{in}}$ and $L_{\text{out}}$ the inner and outer contact polytope, respectively, corresponding to the contact pairs $(v_{i},p_{i})$. Clearly,
$K_{\text{in}}\subset K\subset L\subset L_{\text{out}}.$
Lemma 4.1.
With the above notation and assumptions, we have
$L_{\mathrm{out}}\cap F\subset-d\cdot P_{F}K_{\mathrm{in}}.$
Proof.
Equation (1.1) is equivalent to
(4.2) $P_{F}{\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)}P_{F}=P_{F}{\left(\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}\right)}^{\top}P_{F}=P_{F}.$
[…]
$\sum\limits_{i\in[m]}\upalpha_{i}(1-\left\langle p_{i},x\right\rangle)P_{F}v_{i}\in\sum\limits_{i\in[m]}\upalpha_{i}\cdot P_{F}K_{\text{in}}\cdot(1-\left\langle p_{i},x\right\rangle)\stackrel{{\scriptstyle\left\langle p_{i},x\right\rangle\leq 1}}{{=}}$
[…]
@@ 4.2. Search for a good center
Now, we show that the containment $P_{F}K\subset L\cap F$ guarantees the existence of a translation, after which (4.1) holds.
@@
satisfying identities (1.1), (1.2), and (1.3). Additionally, assume $P_{F}K\subset L\cap F.$
Then there exists a point $z$ in the relative interior of $L\cap F$, contact pairs $\{(v_{1}^{\prime},p_{1}^{\prime}),\dots,(v_{m}^{\prime},p_{m}^{\prime})\}$ of $K-z$ and $L-z$, and weights $c_{1},\dots,c_{m}>0$ such that
${\sum_{i\in[m]}c_{i}p_{i}^{\prime}\otimes v_{i}^{\prime}}=\sum_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i},\quad\text{and}\quad\sum_{i\in[m]}c_{i}P_{F}v_{i}^{\prime}=\sum_{i\in[m]}c_{i}p_{i}^{\prime}=\mathbf{0}.$
@@
This is a purely linear algebraic argument.
We aim at finding a point $z$ in the relative interior of $L\cap F$ such that when we shift the origin to $z$, the shifted contact pairs $(v_{i}-z,p_{i}^{\prime})$ satisfy the required identities for $K-z$ and $L-z$ with appropriate normal vectors $p_{i}^{\prime}$.
First, observe how a contact pair transforms under translation. A pair $(v,p)$ is a contact pair of $K$ and $L$ if and only if, $v$ belongs to the intersection of the boundaries of $K$ and $L$, and the half-space
@@
Set $M=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}.$ Next, we verify the identity $\sum\limits_{i\in[m]}c_{i}(z)p_{i}(z)\otimes v_{i}(z)=M$. Indeed,
$\sum_{i\in[m]}c_{i}(z)p_{i}(z)\otimes v_{i}(z)=\sum_{i\in[m]}\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)\frac{p_{i}}{1-\left\langle z,p_{i}\right\rangle}\otimes(v_{i}-z)=\sum_{i\in[m]}\upalpha_{i}p_{i}\otimes(v_{i}-z)=$
@@
$P_{F}\left(\sum_{i\in[m]}c_{i}(z)v_{i}(z)\right)$ $=\sum_{i\in[m]}\upalpha_{i}\left(1-\left\langle z,p_{i}\right\rangle\right)P_{F}(v_{i}-z)$
@@
Using the identities $\sum\limits_{i\in[m]}\upalpha_{i}p_{i}=\mathbf{0}$ and $M=\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\otimes v_{i}$, we can simplify:
$\sum\limits_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)=\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}-P_{F}\left(\sum\limits_{i\in[m]}\upalpha_{i}v_{i}\otimes p_{i}\right)z-P_{F}z\sum\limits_{i\in[m]}\upalpha_{i}+P_{F}z\left\langle z,\sum\limits_{i\in[m]}\upalpha_{i}p_{i}\right\rangle$
$=\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}-P_{F}{M}^{\top}z-P_{F}z\left(\sum\limits_{i\in[m]}\upalpha_{i}\right)+P_{F}z\cdot 0=\left(\sum\limits_{i\in[m]}\upalpha_{i}P_{F}v_{i}\right)-P_{F}{M}^{\top}z-dP_{F}z.$
Set
$z=P_{F}\left(\frac{\sum\limits_{i\in[m]}\upalpha_{i}v_{i}}{d+1}\right).$
Since by our assumption $P_{F}K\subset L\cap F,$ and since $\sum\limits_{i\in[m]}\upalpha_{i}=d$, and $\mathbf{0}\in\mathrm{int}K$, we conclude that $z$ is in the relative interior of ${L\cap F}$. Thus, $\sum\limits_{i\in[m]}c_{i}(z)P_{F}v_{i}(z)=\mathbf{0}$, as desired. ∎
4.3. Proof of Theorem 1.5
We use Lemma 4.2 to find a “good center” $z.$ The inclusion follows from Lemma 4.1 and the inclusion $K_{\text{in}}\subset K\subset L\subset L_{\text{out}},$ where $K_{\text{in}}$ and $L_{\text{out}}$ are as defined in Subsection 4.1. In the case when $K$ is symmetric with respect to $F,$ the inclusion $P_{F}K\subset L\cap F$ holds.
4.4. Application: right cones
@@
As an application of Lemma 4.1, we obtain the following result, whose proof is essentially the same as that of Theorem 1.5:
Proposition 4.3.
@@
$(K-z)\cap\ell\subseteq d(C_{0}-z)\cap\ell.$
In particular, if $K$ is a convex body on the plane, and $\Delta$ is the largest area isosceles triangle in $K$ with a given direction of its axis of symmetry $\ell$, then with the right choice of $z\in\Delta\cap\ell$, we have $(K-z)\cap\ell\subseteq 2(\Delta-z)\cap\ell$. Note that the factor 2 in this formula is optimal, as shown by the example when $K$ is a square and $\ell$ is parallel to one of its diagonals.
@@ 4.5. No containment in the whole space
Contrasting Theorem 1.5, we show that there is no meaningful inclusion in ${F}^{\bot}$. Surprisingly, the example is provided by ellipsoids.
Proposition 4.4.
Fix an integer $0\leq s\leq d-2$ and a positive real $\lambda$. Let $E$ be an origin-centered ellipsoid in $\mathbb{R}^{d}$ whose principal semi-axes have lengths $\lambda,\lambda^{2},\ldots,\lambda^{d}$. Let $E^{\prime}\in\mathcal{A}_{F}\!\left(\mathbf{B}^{d}\right)$ be the solution to Problem 2, and let the $s$-dimensional subspace $F$ be the axis of $E^{\prime}$. Then
@@
Proposition 4.6.
Let $K$ be a convex body in $\mathbb{R}^{d}$, and let $A:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be the diagonal linear operator with diagonal entries $\lambda,\lambda^{2},\ldots,\lambda^{d}$. Let $K^{\prime}=A^{\prime}K$ be a solution to Problem 4 with some linear operator $A^{\prime}$, for $K$ and $L=AK$, with any $d-s\geq 2$. Then, for any $\mu>0$ there is some value of $\lambda>0$ such that
@@
    for any $\mu^{\prime}<\mu$, no homothetic copy $x+\mu^{\prime}K^{\prime}$ of $K^{\prime}$ contains $L$,
@@
    $\frac{\operatorname{vol}\nolimits_{d}K^{\prime}}{\operatorname{vol}\nolimits_{d}L}\leq\frac{1}{\mu}$.
@@
Let $B_{\mathrm{is}}$ be a largest ball contained in $K$, and let $B_{\mathrm{cs}}$ be the smallest ball containing $K$. Assume that $B_{\mathrm{cs}}=y+\nu B_{\mathrm{is}}$ for $\nu\geq 1$, $y\in\mathbb{R}^{d}$.
Let $\bar{\mu}>0$ be the smallest number such that a homothetic copy of $K^{\prime}$ of ratio $\bar{\mu}$ contains $L$. Then there is a homothetic copy of $A^{\prime}B_{\mathrm{cs}}$ of ratio $\bar{\mu}$ that contains $AB_{\mathrm{is}}$. Consequently, there is a homothetic copy of $A^{\prime}B_{\mathrm{is}}$ of ratio $\bar{\mu}\nu^{2}$ that contains $AB_{\mathrm{cs}}$. Clearly, $A^{\prime}B_{\mathrm{is}}\subseteq L\subseteq AB_{\mathrm{cs}}$. Thus, by the previous example, $\bar{\mu}\nu^{2}\geq\lambda$, implying that $\bar{\mu}\geq\frac{\lambda}{\nu^{2}}$, proving (a).
@@ 5.1. Contact points on the sphere
Let the origin-centered ellipsoid $E=A\mathbf{B}^{d}$ be a solution to Problem 1, that is John’s problem for ellipsoids with a fixed axis of revolution for a convex body $K$ in $\mathbb{R}^{d}$. Then there are unit vectors $u_{1},\dots,u_{m}$, and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that $(Au_{1},{A}^{-1}u_{1}),\dots,(Au_{m},{A}^{-1}u_{m})$ are contact pairs of $E$ and $K$, and
@@
Note that a standard argument using Carathéodory’s theorem for the space $\mathcal{R}$ shows that the number $m$ of points needed in Corollary 5.1 is at most $2d^{2}+d+1$.
@@
Assume that $E=A\mathbf{B}^{d}$ is a solution to Problem 2, that is John’s problem for ellipsoids with any axis of revolution for a convex body $K$ in $\mathbb{R}^{d}$. Then there exist $m\leq 2d^{2}+d+1$ unit vectors $u_{1},\dots,u_{m}$ and positive weights $\upalpha_{1},\dots,\upalpha_{m}$ such that the conclusions of Corollary 5.1 hold as well as
@@
We will call a position $K^{\prime}\in\mathcal{A}_{F}\!\left(K\right)$ with fixed axis $F$ of a convex body $K$ the John position with fixed axis $F$, if the standard unit ball is a solution to John’s general problem with fixed axis $F$ for $K^{\prime}$ and $\mathbf{B}^{d}.$